# open-sci-ref-0.01: open and reproducible reference baselines for language model and dataset comparison

### v2

@@ Abstract
We introduce open-sci-ref, a family of dense transformer models trained as research baselines across multiple model (0.13B to 1.7B parameters) and token scales (up to 1T) on 8 recent open reference datasets. Evaluating the models on various standardized benchmarks, our training runs set establishes reference points that enable researchers to assess the sanity and quality of alternative training approaches across scales and datasets. Intermediate checkpoints allow comparison and studying of the training dynamics. The established reference baselines allow training procedures to be compared through their scaling trends, aligning them on a common compute axis. Comparison of open reference datasets reveals that training on NemoTron-CC HQ consistently outperforms other reference datasets, followed by DCLM-baseline and FineWeb-Edu. In addition to intermediate training checkpoints, the release includes logs, code, and downstream evaluations to simplify reproduction, standardize comparison, and facilitate future research  [^1] Model weights and intermediate checkpoints are available at open-sci-ref-0.01 collection; code for reproducing training, evaluation and raw experiments data open-sci-ref-0.01 repository..
@@ 1 Introduction
To make guided progress in researching and improving foundation models and datasets, it is necessary to compare training procedures across scales. Yet reproducible and strong reference baselines for such comparisons are scarce, with existing ones relying on outdated recipes, older datasets, or being restricted to small scales. In this work, we provide a set of training runs and resulting base language models as improved reference baselines. These can be used to compare and examine other training procedures, assessing their sanity and quality across open reference datasets and scales. We rely on classical dense transformers implementations, as in works like Pythia biderman2023 , DCLM li2025 , OLMo(groeneveld-etal-2024-olmo, ), PolyPythias vanderwal2025 , DataDecide magnusson2025datadecide that use fully open, reproducible training pipelines and also provide intermediate checkpoints for studying training dynamics. Our reference baselines allow comparisons across four models scales (up to 1.7B parameters), trained on eight different datasets and up to larger 1T tokens scale, using a modern training procedure which ensures strong baselines for comparison.
@@ 2.1 Datasets, tokenization, and training details
We consider 8 datasets: 7 established references, namely C4 raffel2020exploring , Pile gao2020pile , SlimPajama cerebras2023slimpajama , FineWeb-Edu-1.4T (v1.0.0) penedo2024 , DCLM-baseline li2025 , Nemotron-CC-HQ su2025nemotroncc , HPLT-2.0 (english subset) burchell2025 , and CommonCorpus langlais2025 , a multi-lingual Common Crawl–based data pool not filtered for training. We tokenize datasets using the GPT-NeoX-20B tokenizer black2022gptneox , resulting in a vocabulary size of 50,304. For all datasets, we train reference dense transformer models with 130M, 400M, 1.3B, and 1.7B parameters using Megatron-LM megatron-lm . Following other strong reference models like Llama, Qwen, DCLM, and DeepSeek, we employ QK normalization naseer2021 in the model architecture for more stable training and constant learning rate with linear cooldown schedule (also known as WSD or trapezoid schedule (hu2024minicpm, ; hagele2024scaling, )), which facilitates training resumption and better supports scaling law derivation across token budgets.
@@ 2.2 Model architecture
We follow a custom design starting from the Llama (touvron2023llamaopenefficientfoundation, ) architecture. Different from standard Llama, we include biases in all linear layers (QKV projections, attention, and FFN), as removing them in ablation experiments at the 1.3B, 50BT scale resulted in performance deterioration, and we employ QK normalization (henry2020qknorm, ) for improved training stability and better performance. We use SwiGLU (shazeer2020glu, ) and tied embedding weights across all model scales. Tab. 1 summarizes the model configurations. Hereafter, we refer to this model design as open-sci-ref.
@@ 2.3 Sanity checks and downstream evaluation
To validate the training procedure of our reference models, we perform a comparison with established baselines on the open datasets C4 and FineWeb-Edu, using open-sourced HuggingFace (HF) models with 1.7B parameters, trained on 350B tokens (denoted as HF-ref). We evaluate both HF-ref and open-sci-ref on standardized benchmarks (Tab. 2). Since open-sci-ref matches or outperforms HF-ref—despite the latter being trained on 50B more tokens—we conclude that our training procedure is well established and tuned for a given 1.7B model scale.
@@
Evaluation. Having establishing the soundness of our training procedure, we execute training of reference baselines and evaluate every intermediate checkpoint using LM-eval-harness eval-harness . We evaluate on Copa copa , Openbookqa mihaylov2018OpenBookQA , Lambada-openai paperno2016lambada , Winogrande sakaguchi2019winogrande on 0-shot setting, MMLU hendrycks2021measuring in 5-shot, and Commonsense-qa talmor2019commonsenseqa , Piqa bisk2020piqa , Arc-challenge clark2018thinkarx , Arc-easy, Hellaswag zellers2019hellaswag , Boolq clark2019boolq with 10-shot, resulting in 11 evaluations in total.
Control experiments. We perform additional experiments, testing the effect of different context lengths (4096, 8192, 16384 tokens) on the 1.7B model trained for 1T tokens, and compare various learning rate schedules (Suppl.  A.3). For additional hyperparameter tuning details, see Suppl. Sec. B
@@ 3.1 Dataset comparison across scales
Fig.˜1 reports the 11-task average performance across four model sizes, each trained on a different dataset at the 300B-token scale. Our results clearly show that recent state-of-the-art datasets such as Nemotron-CC-HQ and DCLM provide better performance, with gaps widening for larger models, while the ranking of top datasets remains stable across scales. By comparison, CommonCorpus, as an example of a multilingual data pool not filtered for base model training, unsurprisingly performs worse than English-dominant reference datasets filtered and curated for training.
Fig.˜2 shows task-level performance for the 1.7B model trained on 300B tokens on the reference datasets. We see pronounced differences between datasets on some evaluation tasks, though the overall ranking remains consistent. On MMLU, dataset comparison is difficult due to the lack of signal for all datasets except Nemotron-CC-HQ. This is likely because this dataset contains generated instruction-like samples with various rephrasings, boosting the ability to deal with multiple choice questions in MMLU, whereas other datasets yield scores close to random guessing (0.25). We additionally report results for the three top-performing datasets at the 1T-tokens scale (see Suppl. Sec. A, Tab. 3). For details on distributed training, see Suppl. Sec. C
3.2 Scaling trends based comparison
Fig.˜3 shows scaling behavior of our open-sci-ref reference baseline models when varying the training compute budget (increasing both model and token scale) for different datasets. By performing measurements across selected reference scales, a consistent dataset ranking can be established by looking at scaling trends across a broad span while aligning evaluated models on a common compute axis. The scaling trends can hint at promising strong datasets even given a low training budget setting and without deriving full scaling laws. Interestingly, our reference baseline 1.7B model trained on Nemotron-cc HQ for 1T token matches SmolLM2-1.7B (both 0.66 average score), achieving competitive results despite being trained with a simple single stage procedure on far fewer tokens (1T vs 11T), thus using roughly $11\times$ less compute (see also Tab. 5).
@@ 3.2 Scaling trends based comparison
In addition to english dominant pre-trained baseline models, we take for comparison multi-lingually pretrained EuroLLM martins2024eurollm ; martins2025eurollm as it represents so far the strongest multi-lingual model coming from EU based efforts (significantly outpeforming similar motivated EU based models like Salamandra, Teuken or Occiglot). Well suited for comparison to our reference baselines, EuroLLM-1.7B used FineWeb-Edu as core part of its pre-training dataset, having at least 50% english data while adding multiple languages to reach total of 4T tokens, making it an interesting case for comparison.
It is important not to take the slopes of the scaling trends as indication how scalable the learning procedure is, without comparing the points to other references. For instance, for its model scale and invested compute, EuroLLM 1.7B strongly underperforms other learning procedures when properly taking into account pre-training compute. Eg, it has same performance level as SmolLM2 360M which has smaller model scale and used substantially less compute as evident from Fig.˜3. It underperforms 1.7B open-sci-ref models despite those using more than 10x less compute (see also Tab. 3). DCLM-1B model that has slightly less compute than EuroLLM 1.7B strongly outperforms it. Such points on smaller scales that are highly suboptimal can create the illusion of a learning procedure with strong scaling trend, as in the case of slope given for EuroLLM-1.7B - EuroLLM-9B line. Comparing EuroLLM-1.7B to other established references on common compute axis helps to avoid the illusion, revealing that steep slope towards EuroLLM-9B is due to the suboptimal performance on smaller 1.7B scale and not due to procedure’s scalability.
Figure 3: Scaling trends: open-sci-ref reference baselines (0.13, 0.4, 1.3, 1.7B model scales and 300B, 1T token scales) using 8 different datasets (solid lines) compared to other models (SmolLM2, Qwen 2.5, EuroLLM, DCLM) across various model and token scales (dashed lines), aligned on common compute axis (all tested models are base models). Evaluating reference baselines that use the same training procedure across scales while accounting for training compute provides dataset comparison and ranking. Nemotron-$C$-HQ, DCLM-baseline and FineWeb-Edu are strong reference datasets leading the ranking.
Tab. 3 shows more details on the per-task and average performance numbers for our reference baselines on 1.7B model and 300B, 1T token scales, providing a comparison to other models of similar sizes and with similar or different token budgets. We observe that EuroLLM-1.7B underperforms across all english benchmarks open-sci-ref-1.7B trained on 1T tokens of FineWeb-Edu data, and even on 300B tokens of FineWeb-Edu, open-sci-ref-1.7B still prevails on many benchmarks despite more than 10x less pre-training compute (more than 6x less compute on English, when comparing 300B vs. 2T tokens). This suggests that mixing languages can substantially weaken English benchmark performance and reduce problem-solving ability, as we see deterioration on English benchmarks despite substantially greater compute spent on English content.
@@ 4 Conclusion & Outlook
Using the established reference baselines, we compare performance across dataset, obtaining dataset rankings that remain consistent across scales (Fig. 1,  3) and over the course of training (Fig. 1,  2). This consistency across scales and training durations provides stronger validation than any single-scale comparison, strengthening our analysis. We further compare against external models, aligning training procedures and reference baselines in a common frame of shared compute axis (Fig. 3). This reveals how weaker or stronger procedures can be identified using small scale baseline experiments. Nevertheless, establishing comparisons that retain validity across broad scales remains an open challenge and will require the derivation of full scaling laws nezhurina2025scaling ; we leave this to follow-up work.
Having derived reference baselines for comparison with any other learning procedures, in future work we will: (a) study permissively licensed datasets—CommonCorpus langlais2025commoncorpuslargestcollection , CommonPile kandpal2025commonpilev018tb , MixtureVitae mixturevitae2025dataset —and compare them to reference datasets which contain non-permissive content; (b) evaluate mixture-of-experts architectures against dense transformers; and (c) derive full scaling laws to enable robust and accurate comparisons at larger scales not covered here.

### v3

@@ Abstract
We introduce open-sci-ref, a family of dense transformer models trained as research baselines across multiple model (0.13B to 1.7B parameters) and token scales (up to 1T) on 8 recent open reference datasets. Evaluating the models on various standardized benchmarks, our training runs set establishes reference points that enable researchers to assess the sanity and quality of alternative training approaches across scales and datasets. Intermediate checkpoints allow comparison and studying of the training dynamics. The established reference baselines allow training procedures to be compared through their scaling trends, aligning them on a common compute axis. Comparison of open reference datasets reveals that training on NemoTron-CC HQ consistently outperforms other reference datasets, followed by DCLM-baseline and FineWeb-Edu. In addition to intermediate training checkpoints, the release includes logs, code, and downstream evaluations to simplify reproduction, standardize comparison, and facilitate future research  [^1] Model weights and intermediate checkpoints are available at https://huggingface.co/collections/open-sci/open-sci-ref-001; code for reproducing training, evaluation and raw experiments data https://github.com/LAION-AI/open-sci-ref-0.01.
@@ 1 Introduction
To make guided progress in researching and improving foundation models and datasets, it is necessary to compare training procedures across scales. Yet reproducible and strong reference baselines for such comparisons are scarce, with existing ones relying on outdated recipes, older datasets, or being restricted to small scales. In this work, we provide a set of training runs and resulting base language models as improved reference baselines. These can be used to compare and examine other training procedures, assessing their sanity and quality across open reference datasets and scales. We rely on classical dense transformers implementations, as in works like Pythia (Biderman et al. 2023), DCLM (Li et al. 2024), OLMo(Groeneveld et al. 2024), PolyPythias (van der Wal et al. 2025), DataDecide (Magnusson et al. 2025) that use fully open, reproducible training pipelines and also provide intermediate checkpoints for studying training dynamics. Our reference baselines allow comparisons across four models scales (up to 1.7B parameters), trained on eight different datasets and up to larger 1T tokens scale, using a modern training procedure which ensures strong baselines for comparison.
@@ 2.1 Datasets, tokenization, and training details
We consider 8 datasets: 7 established references, namely C4 (Raffel et al. 2020), Pile (Gao et al. 2020), SlimPajama (Soboleva et al. 2023), FineWeb-Edu-1.4T (v1.0.0) (Penedo et al. 2024), DCLM-baseline (Li et al. 2024), Nemotron-CC-HQ (Su et al. 2025), HPLT-2.0 (english subset) (Burchell et al. 2025), and CommonCorpus (Langlais et al. 2025), a multi-lingual Common Crawl–based data pool not filtered for training. We tokenize datasets using the GPT-NeoX-20B tokenizer (Black et al. 2022), resulting in a vocabulary size of 50,304. For all datasets, we train reference dense transformer models with 130M, 400M, 1.3B, and 1.7B parameters using Megatron-LM (Shoeybi et al. 2019). Following other strong reference models like Llama, Qwen, DCLM, and DeepSeek, we employ QK-Norm (Henry et al. 2020) in the model architecture for more stable training and constant learning rate with linear cooldown schedule (also known as WSD or trapezoid schedule (Hu et al. 2024; Hägele et al. 2024)), which allows resuming training from intermediate states easily and also saves compute cost when performing experiments across various token budgets, for instance as required for scaling law derivation.
@@ 2.2 Model architecture
We follow a custom design starting from the Llama (Touvron et al. 2023) architecture. Different from standard Llama, we include biases in all linear layers (QKV projections, attention, and FFN), as removing them in ablation experiments at the 1.3B, 50BT scale resulted in performance deterioration, and employ QK-Norm for improved training stability and better performance (Wortsman et al. 2023). We use SwiGLU (Shazeer 2020) and tied embedding weights across all model scales. Tab. 1 summarizes the model configurations. Hereafter, we refer to this model design as open-sci-ref.
@@
2.3 Hyperparameter tuning
Tokens Global bs (tokens) iters lr warmup
50B 2.65M 18839 3e-3 6000
50B 4.03M 12406 6e-3 1000
300B 2.09M 143052 3e-3 5000
Table 2: Hyperparameters for cosine lr schedule
Tokens Global bs (tokens) Iters lr Warmup Cooldown (20%)
50B 2.65M 18839 $2\cdot 10^{-3}$ 6000 3767
50B 4.12M 11921 $4\cdot 10^{-3}$ 1000 2384
300B 2.09M 143052 $1\cdot 10^{-3}$ 5000 28610
300B 4.12M 72661 $4\cdot 10^{-3}$ 25000 14532
1T 4.12M 242204 $4\cdot 10^{-3}$ 25000 48440
We perform hyperparameter tuning by first measuring test loss on C4 and FineWeb-Edu, selecting hyperparameters that result in the lowest test loss on the datasets using 50B training runs with 1.3B model scale.
Learning rate schedules. We first tune hyperparams using standard cosine learning rate schedule (Tab. 2). We execute baseline training on the C4 dataset with 50B and 300B tokens, and obtain hyperparams for cosine learning schedule that correspond to minimal observed test loss. We then switch to constant learning + cooldown (WSD) and tune learning rate until we match or outperform minimal test loss observed with cosine schedule.
After achieving the match and thus checking WSD training procedure sanity, we switch to WSD schedule. The tuned WSD schedule training procedure (Tab. 3) is then used for all follow-up experiments and all model scales (0.13B, 0.4B, 1.3B, 1.7B). Tuned procedure uses weight decay of 0.05, global batch size of 4M tokens, cooldown of 20% of total training duration. Warmup and learning rate depend on the chosen total token budget. For larger token budgets >> 300B, a large warmup of 25 000 iterations turned out to give performance gains.
@@ 2.4 Sanity checks and downstream evaluation
To validate the tuned training procedure of our reference models, we perform a comparison with established baselines on the open datasets C4 and FineWeb-Edu, using open-source HuggingFace (HF) models with 1.7B parameters, trained on 350B tokens (denoted as HF-ref). We evaluate both HF-ref and open-sci-ref on standardized benchmarks (Tab. 4). Since open-sci-ref matches or outperforms HF-ref—despite the latter being trained on 50B more tokens—we conclude that our training procedure is well established and tuned for a given 1.7B model scale.
@@
Evaluation. Having establishing the soundness of our training procedure, we execute training of reference baselines and evaluate every intermediate checkpoint using LM-eval-harness (Gao et al. 2021). We evaluate on Copa (Roemmele, Bejan, and Gordon 2011), Openbookqa (Mihaylov et al. 2018), Lambada-openai (Paperno et al. 2016), Winogrande (Sakaguchi et al. 2021) on 0-shot setting, MMLU (Hendrycks et al. 2021) in 5-shot, and Commonsense-qa (Talmor et al. 2019), Piqa (Bisk et al. 2020), Arc-challenge (Clark et al. 2018), Arc-easy, Hellaswag (Zellers et al. 2019), Boolq (Clark et al. 2019) with 10-shot, resulting in 11 evaluations in total.
Control experiments. We perform additional experiments, testing the effect of different context lengths (4096, 8192, 16384 tokens) on the 1.7B model trained for 1T tokens (Suppl. Sec. A.1).
@@ 3.1 Dataset comparison across scales
Reference baselines on 300B tokens. Fig.˜1 reports the 11-task average performance across four model sizes, each trained on a different dataset at the 300B-token scale. Our results clearly show that recent state-of-the-art datasets such as Nemotron-CC-HQ and DCLM provide better performance, with gaps widening for larger models, while the ranking of top datasets remains stable across scales. By comparison, CommonCorpus, as an example of a multilingual data pool not filtered for base model training, unsurprisingly performs worse than English-dominant reference datasets filtered and curated for training.
Fig.˜2 shows task-level performance for the 1.7B model trained on 300B tokens on the reference datasets. We see pronounced differences between datasets on some evaluation tasks, though the overall ranking remains consistent. On MMLU, dataset comparison is difficult due to the lack of signal for all datasets except Nemotron-CC-HQ. This is likely because this dataset contains generated instruction-like samples with various rephrasings, boosting the ability to deal with multiple choice questions in MMLU, whereas other datasets yield scores close to random guessing (0.25). For details on distributed training, see Suppl. Sec. B
@@
Reference baselines on 1T tokens. Fig.˜3 shows the performance of training the 1.7B model for the top 3 datasets while increasing the budget to 1T tokens. The ranking remains consistent, Nemotron staying mostly on the top, followed or matched by DCLM and FineWeb-Edu (with the exception of Lambada, where DCLM dominates), while performance is going up as expected when increasing token scale.
@@
Tab. 5 shows the per-task and average performance numbers for our reference baselines on 1.7B model and 300B, 1T token scales, providing a comparison to other models of similar sizes and with similar or different token budgets. In addition to strong english dominant pre-trained baseline models like DCLM-1B, we take for comparison multi-lingually pretrained EuroLLM (Martins et al. 2024, 2025) as it represents so far the strongest multi-lingual model coming from EU based efforts (significantly outpeforming similar motivated EU based models like Salamandra, Teuken or Occiglot). Well suited for comparison to our reference baselines, EuroLLM-1.7B used FineWeb-Edu as core part of its pre-training dataset, having at least 50% english data while adding multiple languages to reach total of 4T tokens, making it an interesting case for comparison. We observe that EuroLLM-1.7B underperforms across all english benchmarks open-sci-ref-1.7B trained on 1T tokens of FineWeb-Edu data, and even on 300B tokens of FineWeb-Edu, open-sci-ref-1.7B still prevails on many benchmarks. This again shows that mixing languages can lead to strongly diminished problem solving performance, even if using substantially more compute for English (2x in case of EuroLLM).
@@
Aligning models on the common compute axis for evaluation as done in Fig. 4 and Tab. 5 can provide a good overview where various learning procedures stand relative to each other, which is often difficult if not making compute that went into the training explicit. Again, for EuroLLM 1.7B we see that for its model scale and invested compute, it strongly underperforms other learning procedures. Eg, it has same performance level as SmolLM2 360M which has smaller model scale and used substantially less compute, also matching or underperforming 1.7B open-sci-ref models that used less compute. DCLM-1B model that has slightly less compute than EuroLLM 1.7B strongly outperforms it. It is therefore important to not take the slopes of the scaling trends as indication how scalable the learning procedure is, without comparing the points to other reference procedures. Points on smaller scales that are highly suboptimal can create the illusion of a strongly scalable learning procedure by making the slope very steep. Eg, slope as defined by EuroLLM-1.7B and EuroLLM-9B (Martins et al. 2025) performance is steep due to EuroLLM-1.7B being heavily suboptimal, as becomes apparent by comparison to reference baselines. Proper scaling trend estimation requires the models across scales to be close to compute optimal. Such measurements are executed for scaling law derivation, which identifies models trained close to compute optimal Pareto front and gives thus much more accurate predictions about learning procedure behavior than scaling trends presented here, however requiring more dense measurements and hyperparameter tuning effort.
3.2 Scaling trends based comparison
Figure 4: Scaling trends of open-sci-ref-0.01 baselines: 0.13B (300BT), 0.4B (300BT), 1.3B (300BT), 1.7B (300BT), and 1.7B (1TT) trained across 8 datasets (solid lines). Dashed lines show comparison with external models across parameter and token scales.
Fig.˜4 shows scaling behavior of our open-sci-ref reference baseline models when varying the training compute budget (increasing both model and token scale) for different datasets. By performing measurements across selected reference scales, a consistent dataset ranking can be established by looking at scaling trends across a broad span while aligning evaluated models on a common compute axis. The scaling trends can hint at promising strong datasets even given a low training budget setting and without deriving full scaling laws. Tab. 5 reports the per-task and average performance numbers, providing a comparison of our 1.7B reference baselines with other models of similar sizes and with similar or different token budgets. Interestingly, our reference baseline 1.7B model trained on Nemotron-cc HQ for 1T token matches SmolLM2-1.7B (both 0.66 average score), achieving competitive results despite being trained with a simple single stage procedure on far fewer tokens (1T vs 11T), thus using roughly $11\times$ less compute.
Drawing further on our reference baselines, we compare open-sci-ref-1.7B trained on FineWeb-Edu (300B and 1T tokens) with EuroLLM-1.7B (Martins et al. 2024), aiming to investigate the challenges of multilingual pretraining. EuroLLM-1.7B relied on FineWeb-Edu as the core of its pre-training corpus, containing at least 50% English and extended with multiple languages to reach a total of 4T tokens. We observe that EuroLLM-1.7B underperforms open-sci-ref-1.7B trained on 1T tokens of FineWeb-Edu across all English benchmarks; notably, even the 300B-token open-sci-ref-1.7B surpasses EuroLLM-1.7B on many tasks. This suggests that mixing languages can substantially weaken English benchmark performance and reduce problem-solving ability, despite EuroLLM’s much larger overall token budget (4T) and substantially greater compute spent on English content ($2\times$) compared with open-sci-ref-1.7B at 1T.
3.3 Further insights from multilingual benchmarks
We perform comparisons on two parallel multilingual benchmarks: MMMLU, which contains translated MMLU instructions by OpenAI and Belebele (Bandarkar et al. 2024).
In Fig.˜6, we show the performance on Belebele while training a 1.7B model on Nemotron-CC-HQ, which echoes the surprising results seen on MMMLU. While the dataset is supposed to be English only, the model trained exhibits above random-chance performance (25%) on languages such as French, Spanish, Danish, or Portuguese. This is a surprising result which ought to be investigated. We hypothesize that it could come from the fact that the tolerance threshold used when filtering for English text lets further substantial data in another languages to slip into Nemontron dataset (Su et al. 2025).
@@ 4 Conclusion & Outlook
Using the established reference baselines, we compare performance across dataset, obtaining dataset rankings that remain consistent across scales (Fig. 1, 4) and over the course of training (Fig. 1, 2). This consistency across scales and training durations provides stronger validation than any single-scale comparison, strengthening our analysis. We further compare against external models, aligning training procedures and reference baselines in a common frame of shared compute axis (Fig. 4). This reveals how weaker or stronger procedures can be identified using small scale baseline experiments. Nevertheless, establishing comparisons that retain validity across broad scales remains an open challenge and will require the derivation of full scaling laws (Nezhurina et al. 2025); we leave this to follow-up work.
Having derived reference baselines for comparison with any other learning procedures, in future work we will: (a) study permissively licensed datasets—CommonCorpus (Langlais et al. 2025), CommonPile (Kandpal et al. 2025), MixtureVitae (Nguyen et al. 2025)—and compare them to reference datasets which contain non-permissive content; (b) evaluate mixture-of-experts architectures against dense transformers; and (c) derive full scaling laws to enable robust and accurate comparisons at larger scales not covered here.